TOWARDS THE FIRST MEASUREMENT OF PARITY VIOLATION IN CHIRAL MOLECULES NEW ATTEMPTS AND FUTURE PROSPECTIVE

\underline{PETER SCHWERDTFEGER}, Centre for Theoretical Chemistry and Physics, The New Zealand Institute for Advanced Study, Massey University Auckland, Auckland, New Zealand.

Parity violation (PV) effects in atomic transitions have been measured and calculated to high accuracy confirming the so-called standard model in particle physics. The Z-boson exchange between electrons and nucleons leads to a small energy difference between enantiomers of chiral molecules. There is, however, no experimental verification yet of this distinct symmetry breaking effect despite many attempts. Current high-resolution optical spectroscopy experiments carried out in the CO_{2} laser frequency range (878-1108 cm-1) in Christian Chardonnets group in Paris achieve resolutions below 1 Hz. Recent calculations in our group applying the standard model show that PV effects in vibrational transitions of chiral methane derivates CFXYZ (X,Y,Z= H, Cl, Br, I) are in the mHz range and below the detection limit. Our research group is therefore searching for better molecules including heavy elements (because of the PV Z^{5} scaling with nuclear charge Z) to achieve enhanced PV splittings in the Hz range. New promising candidates are presented in collaboration with the French PV initiative, which aims at a 100 mHz resolution in Ramsay-Fringes experiments using quantum cascade lasers. Another future alternative is single-molecule spectroscopy in traps at ultra-cold temperatures.